Find the dimensions of the rectangular field of maximum area which can be enclosed with 400 feet of fence.
Answer by nerdybill(7384)

Find the dimensions of the rectangular field of maximum area which can be enclosed with 400 feet of fence.
.
Let w = width of field
then
(400-2w)/2 = length of field
(200-w) = length of field
.
area = w(200-w)
area = 200w-w^2
area = -w^2+200w
.
Looking at the coefficient for the w^2 term, we see that it is negative. This indicates that the parabola opens downward and finding the vertex will give you the "maximum".
.
Standard vertex form is:
y= a(x-h)^2+k
where
(h,k) is the vertex
.
Convert our equation into that form by "completing the square":
area = -w^2+200w
area = -(w^2-200w)
area = -(w^2-200w+10000) + 10000
area = -(w-100)^2 + 10000
.
From the above, we see that the vertex is:
(h,k) = (100, 10000)
.
This says that when the width=100 feet, the area will be maximized at 10000 square feet.
.
Solution:
width = 100 feet
length = 200-w = 200-100 = 100 feet